OpenStudy (anonymous):

Sam scored 59, 64, and 68 on three tests. How many points should he get on the fourth test to get an average score of 70? 63.67 points 65.25 points 75 points 89 points

Mehek (mehek14):

what is the average of the first 3 tests?

OpenStudy (anonymous):

x1+x2+x3+x4 all over number of tests = ave

OpenStudy (anonymous):

59+64+68+x all over 4 = 70 that would be your equation

OpenStudy (anonymous):

\[\frac{1}{4} (x+68+64+59)=70 \] x = 89